# Efficient Feature Fusion for UAV Object Detection

### v1

@@ Abstract
Object detection in unmanned aerial vehicle (UAV) remote sensing images poses significant challenges due to unstable image quality, small object sizes, complex backgrounds, and environmental occlusions. Small objects, in particular, occupy minimal portions of images, making their accurate detection highly difficult. Existing multi-scale feature fusion methods address these challenges to some extent by aggregating features across different resolutions. However, these methods often fail to effectively balance classification and localization performance for small objects, primarily due to insufficient feature representation and imbalanced network information flow. In this paper, we propose a novel feature fusion framework specifically designed for UAV object detection tasks to enhance both localization accuracy and classification performance. The proposed framework integrates hybrid upsampling and downsampling modules, enabling feature maps from different network depths to be flexibly adjusted to arbitrary resolutions. This design facilitates cross-layer connections and multi-scale feature fusion, ensuring improved representation of small objects. Our approach leverages hybrid downsampling to enhance fine-grained feature representation, improving spatial localization of small targets, even under complex conditions. Simultaneously, the upsampling module aggregates global contextual information, optimizing feature consistency across scales and enhancing classification robustness in cluttered scenes. Experimental results on two public UAV datasets demonstrate the effectiveness of the proposed framework. Integrated into the YOLO-V10 model, our method achieves a 2% improvement in average precision (AP) compared to the baseline YOLO-V10 model, while maintaining the same number of parameters. These results highlight the potential of our framework for accurate and efficient UAV object detection.
@@ I Introduction
Unlike traditional ground-based object detection, UAV-based object detection presents unique challenges. The aerial perspective introduces unstable image quality, small object sizes, complex backgrounds, and frequent environmental occlusions. Small objects, in particular, occupy only a minimal fraction of the image, further complicating accurate detection. These challenges necessitate innovative solutions to address the limitations of existing object detection frameworks.
@@
Figure 1: llustration of our framework. (a) Backbone with fusion down sample (FDS) module added for collecting shallow layer features. (b) Neck of network. (c) Fusion up sample (FUS) module for collecting deeper layer features, (d) Detection head (e) Fusion multi-head self-attention (FMSA) module for the feature fusion of output from (a), (b), and (c).
@@
Currently, deep learning techniques are the dominant approach for UAV-based object detection [51, 52, 53, 54, 55, 56]. Existing methods can be categorized into three main groups. The first group includes two-stage detectors, such as Faster R-CNN, which use a region proposal network (RPN) to generate candidate regions, followed by classification and regression. These models are known for their high accuracy but often suffer from slower inference speeds, limiting their applicability in real-time scenarios. The second group comprises single-stage detectors, such as the YOLO series, which directly predict object categories and bounding boxes in a single step. While these models offer faster inference, they face challenges in detecting small objects due to insufficient feature representation. The third group consists of transformer-based detectors, such as Vision Transformer (ViT) and Detection Transformer (DETR), which excel at modeling global features and large receptive fields. However, their computational complexity increases significantly when processing high-resolution UAV images, and their performance on small object detection often falls short compared to CNN-based methods.
@@
Object detection [58, 59, 60, 61, 62] has recently seen advancements driven by diffusion models [36, 29, 63, 64], which have demonstrated remarkable capabilities in generating high-resolution synthetic data and enhancing model robustness in challenging scenarios. These models are widely applied in domains such as autonomous driving, medical imaging, and remote sensing, where precise detection and localization are critical. Unlike classification tasks, object detection requires accurate classification information alongside precise localization of objects. One prominent challenge in mainstream CNN-based object detection is the imbalance between classification and localization information. As network architectures become deeper, the enhancement of semantic features often leads to a degradation of spatial details. Additionally, detection heads in multi-scale architectures, such as YOLO-V10, are limited in their ability to balance high-resolution and low-resolution features, resulting in redundant structures and reduced localization accuracy, particularly for small objects.
@@
To address these challenges, we propose a novel framework for feature fusion that integrates cross-layer connections and multi-scale fusion. As illustrated in Fig. 1, our method seamlessly integrates into existing CNN-based architectures. By leveraging Fusion Down Sampling (FDS) and Fusion Up Sampling (FUS) modules, it captures valuable features from shallow and deep network layers. These features are fused with the network’s output using a Fusion Multi-Head Self-Attention (FMSA) module, significantly enhancing detection performance in UAV scenarios. Specifically, the FDS module enhances fine-grained feature representation, which is crucial for improving the spatial localization of small targets, even under complex conditions. Meanwhile, the FUS module aggregates global contextual information, optimizing feature consistency across scales and enhancing classification robustness in cluttered scenes. This dual approach enhances both localization accuracy and classification performance. Our framework is designed to be flexible and can be easily integrated into various CNN architectures, making it a comprehensive enhancement for UAV object detection tasks.
@@ II-A Traditional Object Detection Methods
Traditional object detection algorithms rely on handcrafted feature extraction and sliding window techniques. These methods typically involve three key stages: region proposal, feature extraction, and classification regression. During the region proposal stage, potential object locations are identified. Handcrafted methods are then employed to extract features from these candidate regions, which are subsequently classified using traditional classifiers. Representative approaches include the Viola–Jones detector [2] and the HOG pedestrian detector [3]. While these methods laid the groundwork for object detection, they suffer from high computational complexity, limited feature representation capabilities, and optimization difficulties, making them unsuitable for modern UAV detection tasks.
Figure 2: llustration of our model architecture. The proposed method is a supplementary component and integrated to the state-of-the-art model YOLO-V10, including FDS, FUS, FMSA modules.
@@ II-B Deep Learning-Based Object Detection
Deep learning has revolutionized object detection by leveraging convolutional neural networks (CNNs). These approaches can be broadly categorized into two-stage and single-stage detection methods.
@@
II-C Transformer-Based Methods and UAV Object Detection
Recent advances in Transformer-based object detection have introduced powerful architectures for global feature modeling. Vision Transformer (ViT) [25, 72, 73] applies the Transformer architecture to computer vision, offering larger receptive fields and flexible weight-sharing strategies. Swin Transformer [26] employs a hierarchical structure with a sliding window approach, while DETR [28] frames object detection as a set prediction task, enabling end-to-end detection. However, these models face challenges in handling high-resolution UAV images due to excessive computational overhead and suboptimal performance for small objects. UAV object detection presents additional challenges, such as small object sizes, complex backgrounds, and environmental occlusions. Multi-scale feature learning [8, 50] has proven effective for addressing these issues. Hybrid approaches combining CNNs and Transformers, such as Conformer [29, 30] and LPSW [32], leverage local feature extraction with CNNs and global feature modeling with Transformers. Other methods, including DIAG-TR [33] and NeXtFormer [71], enhance the representation of small objects by integrating convolutional layers with Transformer blocks.
@@ III Proposed Method
This section introduces our proposed fusion multi-head self-attention (FMSA) framework, designed to address the challenges of UAV object detection by enhancing the balance between classification and localization information, particularly for small-object detection. The FMSA framework integrates two auxiliary modules—fusion down sample (FDS) and fusion up sample (FUS)—to perform multi-scale feature fusion and long-range cross-layer connections.
@@ III-A Overall Architecture
Figure 3: llustration of FMSA module. It is an additional module to CNN-based network. Apart from the main output of the network, the FDS module collects shallow layer features, and the FUS module collects deeper layer features for feature fusion conducted by the FMSA module.
@@ III-B Fusion Multi-Head Self-Attention (FMSA)
Object detection tasks require effective processing of both classification and localization information, but deep networks often experience an imbalance between these two aspects. This issue is exacerbated in small-object detection, where semantic information from deeper layers conflicts with spatial information from shallower layers.
@@
The FMSA module consists of multi-head self-attention (MSA) blocks, a multi-layer perceptron (MLP), and residual connections. Positional information is retained through standard sinusoidal position embeddings. The input feature map $x$ is reshaped into $H\times W$ tokens with dimension $C$, represented as:
$x_{p},\quad p\in\{1,2,\dots,H\times W\}.$
@@
$Z_{0}=\left[x_{1};x_{2};\dots;x_{p}\right]+E_{pos},\quad E_{pos}\in\mathbb{R}^{(H\times W)\times C},$
where $E_{pos}$ denotes positional embeddings. Subsequent layers are computed as:
$Z_{L}^{\prime}=\operatorname{MSA}(\operatorname{LN}(Z_{L-1}))+Z_{L-1},$
$Z_{L}=\operatorname{MLP}(\operatorname{LN}(Z_{L}^{\prime}))+Z_{L}^{\prime}.$
@@
$y=\operatorname{C2F}(\operatorname{MHSA}(FDS_{out}+FUS_{out}+x)+x).$
@@ III-C Fusion Down Sample (FDS)
Figure 4: llustration of FDS module. On the right is the network structure of the FDS module, and on the left is the downsampling module of YOLO-V10 for comparison.
The LADS module applies self-attention to 2×2 patches, aggregates outputs through an MLP, and reshapes them to a reduced resolution. For input feature map $x$:
@@
$\quad Z_{local^{\prime}}\in\mathbb{R}^{\text{(stride}\times\text{stride)}\times C}.$
$Z_{local}=\operatorname{MLP}(\operatorname{MSA}(\operatorname{LN}(Z_{local^{\prime}}))),\quad Z_{local}\in\mathbb{R}^{1\times C}.$
@@
$y=\operatorname{C2F}(\operatorname{Concat}(\operatorname{Conv}(x),Z)).$
@@ III-D Fusion Up Sample (FUS)
Figure 5: llustration of FUS module. It performs up-sampling operation for the target deeper layers.
Feature maps from resolutions $H/32\times W/32$ and $H/16\times W/16$ are processed through global attention upsample (GAUS) blocks to produce $H/8\times W/8$ feature maps. The GAUS operation is defined as:
$Z=\operatorname{Replicate}(\operatorname{MLP}(\operatorname{MSA}(\operatorname{LN}(x)))),$
$\quad x\in\mathbb{R}^{(H\times W)\times C}.$
$Z\in\mathbb{R}^{(H\times W\times\text{stride}\times\text{stride})\times\frac{C}{\text{stride}}}.$
The transformed feature maps are then fused with the FDS output and original network features in the FMSA module, improving detection accuracy by enhancing the distinction between objects and background, as well as handling occlusions effectively.
@@ IV Experiment and Analysis
In the experiment, two public datasets, the VisDrone2019 dataset and the DOTA1.5 dataset, were used to train and test the proposed method.
@@ IV-A Datasets
DOTA1.5: DOTA is a large-scale dataset for object detection in aerial imagery. It is designed for the development and evaluation of object detectors in aerial images, featuring objects with various scales, orientations, and shapes. DOTA-v1.5 consists of 2,806 high-resolution aerial images from diverse scenes and environments, with annotations also provided for extremely small instances (smaller than 10 pixels). Of these, 1,411 images are used as the training set and 458 images as the validation set. The dataset includes 15 common object categories, such as large vehicles, small vehicles, ships, bridges, and airplanes.
@@ IV-B Evaluation Metrics
In object detection metrics, most references for small object detection in UAV-based images use mean average precision (mAP) as the evaluation metric [43, 44, 37]. Here, average precision (AP) is a numerical measure that summarizes the shape of the Precision-Recall curve, used to determine the optimal balance between precision and recall in practice. Therefore, in the comparative experiments, mAP is used as the primary metric, as it provides a comprehensive assessment of the model’s performance. The definition of mAP is as follows:
$mAP=\frac{1}{n}\sum_{i=1}^{n}AP_{i}$
Here, $n$ represents the number of identified object categories, and $AP_{i}$ denotes the area under the Precision-Recall curve for the i-th object category.
@@ IV-C Implementation Details
In this paper, we conduct experiments on the Ubuntu 22.04 system with two NVIDIA GeForce RTX 3090 GPUs, each equipped with 24GB of memory. The experimental environment includes PyTorch 2.0.1, CUDA 11.7, and Python 3.9. We evaluate our method on the VisDrone2019 and DOTAv1.5 datasets, which contain 8k training images and 1k validation images in total. During training, the input feature map dimensions are set to 640 × 640, and the model is trained for a maximum of 1000 epochs. The batch size is 6, and we use 8 workers for data loading. YOLO-V10 is used as the baseline model, with our modules applied to it. The model is trained using the SGD optimizer with a momentum of 0.937, and the initial learning rate is set to 0.01. We adopt the default training configuration of YOLO-V10 with the specific parameter settings as shown in TABLE I.

### v3

@@ Abstract
Object detection in unmanned aerial vehicle (UAV) remote sensing images poses significant challenges due to unstable image quality, small object sizes, complex backgrounds, and environmental occlusions. Small objects, in particular, occupy small portions of images, making their accurate detection highly difficult. Existing multi-scale feature fusion methods address these challenges to some extent by aggregating features across different resolutions. However, they often fail to effectively balance the classification and localization performance for small objects, primarily due to insufficient feature representation and imbalanced network information flow. In this paper, we propose a novel feature fusion framework specifically designed for UAV object detection tasks to enhance both localization accuracy and classification performance. The proposed framework integrates hybrid upsampling and downsampling modules, enabling feature maps from different network depths to be flexibly adjusted to arbitrary resolutions. This design facilitates cross-layer connections and multi-scale feature fusion, ensuring improved representation of small objects. Our approach leverages hybrid downsampling to enhance fine-grained feature representation, improving spatial localization of small targets, even under complex conditions. Simultaneously, the upsampling module aggregates global contextual information, optimizing feature consistency across scales and enhancing classification robustness in cluttered scenes. Experimental results on two public UAV datasets demonstrate the effectiveness of the proposed framework. Integrated into the YOLO-v10 model, our method achieves a 2 percentage points improvement in average precision (AP) compared to the baseline YOLO-v10 model, while maintaining the same number of parameters. These results highlight the potential of our framework for accurate and efficient UAV object detection.
@@ I Introduction
UAV-based object detection presents unique challenges compared with traditional ground-based object detection. The aerial perspective introduces unstable image quality, small object sizes, complex backgrounds, and frequent environmental occlusions. Small objects, in particular, occupy only a small fraction of the image, further complicating accurate detection. These challenges necessitate innovative solutions to address the limitations of existing object detection frameworks.
@@
Figure 1: Illustration of our framework. (a) Backbone with fusion down sample (FDS) module added for collecting shallow layer features. (b) Neck of network. (c) Fusion up sample (FUS) module for collecting deeper layer features. (d) Detection head. (e) Fusion multi-head self-attention (FMSA) module for the feature fusion of outputs from (a), (b), and (c).
@@
Currently, deep learning techniques are the dominant approach for UAV-based object detection [51, 52, 53, 54, 55]. Existing methods can be categorized into three main groups. The first group includes two-stage detectors, such as Faster R-CNN, which use a region proposal network (RPN) to generate candidate regions, followed by classification and regression. These models are known for their high accuracy but often suffer from slower inference speeds, limiting their applicability in real-time scenarios. The second group comprises single-stage detectors, such as the YOLO series, which directly predict object categories and bounding boxes in a single step. While these models offer faster inference, they face challenges in detecting small objects due to insufficient feature representation. The third group consists of Transformer-based detectors, such as Vision Transformer (ViT) and Detection Transformer (DETR), which excel at modeling global features and large receptive fields. However, their computational complexity increases significantly when processing high-resolution UAV images, and their performance on small object detection often falls short compared to CNN-based methods. Object detection [58, 61, 62] has recently seen advancements driven by diffusion models [36, 29, 63, 64], which have demonstrated remarkable capabilities in generating high-resolution synthetic data and enhancing model robustness in challenging scenarios. These models are widely applied in domains such as autonomous driving, medical imaging, and remote sensing, where precise detection and localization are critical.
@@
One prominent challenge for object detection algorithms, e.g., mainstream CNN-based ones, is the imbalance between the classification and localization information. As network architectures become deeper, the enhancement of semantic features often leads to a degradation of spatial details. Additionally, detection heads in multi-scale architectures, such as YOLO-v10, are limited in their ability to balance high-resolution and low-resolution features, resulting in redundant structures and reduced localization accuracy, particularly for small objects.
@@
To address these challenges, we propose a novel framework for feature fusion that integrates cross-layer connections and multi-scale fusion. As illustrated in Fig. 1, our method seamlessly integrates into existing CNN-based architectures. By leveraging Fusion Down Sampling (FDS) and Fusion Up Sampling (FUS) modules, it captures valuable features from shallow and deep network layers. These features are fused with the network’s output using a Fusion Multi-Head Self-Attention (FMSA) module, significantly enhancing the detection performance in UAV scenarios. Specifically, the FDS module enhances fine-grained feature representation, which is crucial for improving the spatial localization of small targets, even under complex conditions. Meanwhile, the FUS module aggregates global contextual information, optimizing feature consistency across scales and enhancing classification robustness in cluttered scenes. This dual approach improves both localization accuracy and classification performance. Our framework is designed to be flexible and can be easily integrated into various CNN architectures, making it a comprehensive enhancement for UAV object detection tasks.
@@ II-A Traditional Object Detection Methods
Traditional object detection algorithms rely on handcrafted feature extraction and sliding window techniques. These methods typically involve three key stages: region proposal, feature extraction, and classification regression. During the region proposal stage, potential object locations are identified. Then, handcrafted methods are employed to extract features from these candidate regions, followed by classification using traditional classifiers. Representative approaches include the Viola–Jones detector [2] and the Histograms of oriented gradients (HOGs) for human detection [3]. While these methods laid the groundwork for object detection, they suffer from high computational complexity, limited feature representation capabilities, and optimization difficulties, making them unsuitable for modern UAV detection tasks.
Figure 2: Illustration of our model architecture. The proposed method is a supplementary component and integrated to the state-of-the-art model YOLO-v10, including FDS, FUS, FMSA modules.
@@ II-B Deep Learning-Based Object Detection
Deep learning has revolutionized object detection by leveraging CNNs. These approaches can be broadly categorized into two-stage and single-stage detection methods.
@@
Recent advances in Transformer-based object detection have introduced powerful architectures for global feature modeling. Vision Transformer (ViT) [25, 72, 73] applies the Transformer architecture to computer vision, offering larger receptive fields and flexible weight-sharing strategies. Swin Transformer [26] employs a hierarchical structure with a sliding window approach, while DETR [28] frames object detection as a set prediction task, enabling end-to-end detection.
II-C UAV Object Detection
While Transformer-based detectors have shown promising global modeling capabilities, their direct application to UAV imagery remains challenging. Specifically, these models face challenges in handling high-resolution UAV images due to excessive computational overhead and suboptimal performance for small objects. UAV object detection presents additional challenges, such as small object sizes, complex backgrounds, and environmental occlusions. Multi-scale feature learning [8, 50] has proven effective for addressing these issues. Hybrid approaches combining CNNs and Transformers, such as Conformer [29] and LPSW [32], leverage local feature extraction with CNNs and global feature modeling with Transformers. Other methods, including DIAG-TR [33] and NeXtFormer [71], enhance the representation of small objects by integrating convolutional layers with Transformer blocks.
@@ III Proposed Method
This section introduces our proposed fusion multi-head self-attention (FMSA) framework, designed to address the challenges of UAV object detection by enhancing the balance between classification and localization information, particularly for small object detection. The FMSA framework integrates two auxiliary modules, FDS and FUS, to perform multi-scale feature fusion and long-range cross-layer connections.
@@ III-A Overall Architecture
Figure 3: Illustration of FMSA module. It is an additional module to CNN-based network. Apart from the main output of the network, the FDS module collects shallow layer features, and the FUS module collects deeper layer features for feature fusion conducted by the FMSA module.
@@ III-B Fusion Multi-Head Self-Attention (FMSA)
Object detection tasks require effective processing of both classification and localization information, but deep networks often experience an imbalance between these two aspects. This issue is exacerbated in small-object detection, where semantic information from deeper layers cannot be effectively balanced with spatial information from shallower layers.
@@
As illustrated in Fig. 3, the FMSA module consists of multi-head self-attention (MSA) blocks, a multi-layer perceptron (MLP), and residual connections. Positional information is retained through standard sinusoidal position embeddings. The input feature map $\bm{x}$ is reshaped into $H\times W$ tokens with dimension $C$, represented as:
$\bm{x}_{p},\quad p\in\{1,2,\dots,H\times W\}.$
@@
$\bm{Z}_{0}=\left[\bm{x}_{1};\bm{x}_{2};\dots;\bm{x}_{p}\right]+\bm{E}_{pos},\quad\bm{E}_{pos}\in\mathbb{R}^{(H\times W)\times C},$
where $\bm{E}_{pos}$ denotes positional embeddings. Subsequent layers are computed as:
$\bm{Z}_{L}^{\prime}=\operatorname{MSA}(\operatorname{LN}(\bm{Z}_{L-1}))+\bm{Z}_{L-1},$
$\bm{Z}_{L}=\operatorname{MLP}(\operatorname{LN}(\bm{Z}_{L}^{\prime}))+\bm{Z}_{L}^{\prime}.$
@@
$\bm{y}=\operatorname{C2F}(\operatorname{MHSA}(FDS_{out}+FUS_{out}+\bm{x})+\bm{x}),$
where C2F is a YOLO-v10’s module for effective feature extraction and fusion.
@@ III-C Fusion Down Sample (FDS)
Figure 4: Illustration of FDS module. On the right is the network structure of the FDS module, and on the left is the downsampling module of YOLO-v10 for comparison.
As shown in Fig. 4, the LADS module applies self-attention to 2×2 patches, aggregates outputs through an MLP, and reshapes them to a reduced resolution. For input feature map $\bm{x}$:
@@
$\quad\bm{Z}_{local^{\prime}}\in\mathbb{R}^{\text{(stride}\times\text{stride)}\times C}.$
$\bm{Z}_{local}=\operatorname{MLP}(\operatorname{MSA}(\operatorname{LN}(\bm{Z}_{local^{\prime}}))),\quad\bm{Z}_{local}\in\mathbb{R}^{1\times C}.$
@@
$\bm{y}=\operatorname{C2F}(\operatorname{Concat}(\operatorname{Conv}(\bm{x}),\bm{Z})).$
@@ III-D Fusion Up Sample (FUS)
Figure 5: Illustration of FUS module. It performs up-sampling operation for the target deeper layers.
As shown in Fig. 5, feature maps from resolutions $H/32\times W/32$ and $H/16\times W/16$ are processed through global attention upsample (GAUS) blocks to produce $H/8\times W/8$ feature maps. The GAUS operation is defined as:
$\bm{Z}=\operatorname{Replicate}(\operatorname{MLP}(\operatorname{MSA}(\operatorname{LN}(\bm{x})))),$
where $\bm{x}\in\mathbb{R}^{(H\times W)\times C}$, and $\bm{Z}\in\mathbb{R}^{(H\times W\times\text{stride}\times\text{stride})\times\frac{C}{\text{stride}}}$.
The transformed feature maps are then fused with the FDS output and the original network features in the FMSA module, improving detection accuracy by enhancing the distinction between objects and background, as well as handling occlusions effectively.
@@ IV Experiment and Analysis
In the experiment, two public datasets, the VisDrone2019 dataset and the DOTA-v1.5 dataset, were used to train and test the proposed method.
@@ IV-A Datasets
DOTA-v1.5: DOTA is a large-scale dataset for object detection in aerial imagery. It is designed for the development and evaluation of object detectors in aerial images, featuring objects with various scales, orientations, and shapes. DOTA-v1.5 consists of 2,806 high-resolution aerial images from diverse scenes and environments, with annotations also provided for extremely small instances (smaller than 10 pixels). Of these, 1,411 images are used as the training set and 458 images as the validation set. The dataset includes 15 common object categories, such as large vehicles, small vehicles, ships, bridges, and airplanes.
@@ IV-B Evaluation Metrics
In object detection metrics, most references for small object detection in UAV-based images use mean average precision (mAP) as the evaluation metric [43, 44, 37]. Here average precision (AP) is a numerical measure that summarizes the shape of the Precision-Recall curve, used to determine the optimal balance between precision and recall in practice. Therefore, in the comparative experiments, mAP is used as the primary metric, as it provides a comprehensive assessment of the model’s performance. The definition of mAP is as follows:
$mAP=\frac{1}{n}\sum_{i=1}^{n}AP_{i}.$
Here $n$ represents the number of identified object categories, and $AP_{i}$ denotes the area under the Precision-Recall curve for the $i$-th object category.
@@ IV-C Implementation Details
In this paper, we conduct experiments on the Ubuntu 22.04 system with two NVIDIA GeForce RTX 3090 GPUs, each equipped with 24GB of memory. The experimental environment includes PyTorch 2.0.1, CUDA 11.7, and Python 3.9. We evaluate our method on the VisDrone2019 and DOTA-v1.5 datasets, which contain 8k training images and 1k validation images in total. During training, the input feature map dimensions are set to 640 × 640, and the model is trained for a maximum of 1000 epochs. The batch size is 6, and we use 8 workers for data loading. YOLO-v10 is used as the baseline model, with our modules applied to it. The model is trained using the SGD optimizer with a momentum of 0.937, and the initial learning rate is set to 0.01. We adopt the default training configuration of YOLO-v10 with the specific parameter settings as shown in TABLE I.